# GnnX-Bench: Unravelling the Utility of Perturbation-based Gnn Explainers through In-depth Benchmarking

### v1

@@ A.7 Feasibility for Factual Explanations
The feasibility metric is commonly used in the context of counterfactual graph explainers because it measures how feasible it is to achieve a specific counterfactual outcome. In other words, it assesses the likelihood of a counterfactual scenario being realized given the constraints and assumptions of the underlying base model. On the other hand, factual explainers aim to explain why a model makes a certain prediction based on the actual input data. They do not involve any counterfactual scenarios, so the feasibility metric is not relevant in this context. Instead, factual explainers may use other metrics such as sufficiency and reproducibility to provide insights into how the model is making its predictions. Therefore, we have not used feasibility metrics for factual explanations.
B Stability
Stability against feature perturbation: To determine the perturbation amount, we apply the following procedure. For each feature we compute its standard deviation . Then we sample a value uniformly at random from . The feature value is perturbed to .
Figure F(left inset) illustrates the outcomes, demonstrating a continuation of the previously observed trends. Among these trends, RCExplainer exhibits the highest stability, whereas GNNExplainer displays the lowest stability. This discrepancy in stability could potentially be ascribed to the transductive nature of GNNExplainer, as opposed to the remaining methods, which are all inductive. Transductive methods do not generalize to unseen data and hence when the topology is perturbed, stability suffers. TAGExplainer is task agnostic and hence shows more more instability compared to the other inductive methods.
@@ B Stability
Adversarial attack on topology: For topology adversarial attack, we follow the flip edge method from [49] with a query size of one. Figure F demonstrates the performance of four factual methods on these evasion attacks. The behavior of the factual methods is similar to the topological noise attack explained in Section 4.2 and the feature perturbations results. When adversarial attack is compared to random perturbations (Fig. 3), we observe higher deterioration in stability, which is expected since adversarial attacks are designed to perturb edges that has higher impact on Gnn output.
B.1 Counterfactual Explainers
@@
Figure G presents the average Jaccard similarity results, indicating the similarity between the counterfactual graph predicted as an explanation for the original graph and the noisy graphs at varying levels of perturbations. Additionally, Figure H demonstrates the performance of different explainers in terms of sufficiency and size as the degree of noise increases. This provides insights into how these explainers handle higher levels of noise.
RCExplainer outperforms other baselines by a significant margin in terms of size and sufficiency across datasets, as shown in Fig. H. However, the Jaccard similarity between RCExplainer and CF for counterfactual graphs is nearly identical, as shown in Fig. G. CF benefits from its transductive training on noisy graphs. CLEAR’s results are not shown for Proteins and Mutagenicity datasets due to scalability issues. In the case of IMDB-B dataset, CLEAR is highly unstable in predicting counterfactual graphs, indicated by a low Jaccard index (Fig. G). Additionally, CLEAR demonstrates high sufficiency but requires a large number of edits, indicating difficulty in finding minimal-edit counterfactuals (Fig. H).
Overall, RCExplainer seems to be the model of choice when topological noise is introduced, and it is significantly faster than because it is inductive. Further, it is better than CLEAR as the latter does not scale for larger datasets and is inferior in terms of sufficiency and size as well.
Stability against explainer instances: In Table 8, we provide an overview of the stability using Jaccard index exhibited among explainer instances trained using three distinct seeds. Similarly, as additional results, here we present the variations in sufficiency (Table M) and the explanation size (Table N) produced by the methods using three different seeds. In terms of sufficiency, the methods show stability while varying the seeds. However, the results are drastically different for explanation size. Table N present the results. We observe that RCExplainer is consistently the most stable method while Cf is worse. The worst stability is shown by Clear and this observation is consistent with the previous results.
@@
Stability against Gnn architectures: Table O shows the stability of the explainers across different Gnn architectures. Similar to our factual setting (Table 7), we assess the stability by computing the Jaccard coefficient between the explained predictions of the indicated Gnn architecture and the default Gcn model. Unsurprisingly, the stability of the explainer highly depends on the dataset. RCExplainer is also the most stable among all the explainers, and the produced high values indicate that the method is agnostic towards the variations in different message aggregating schemes of the architectures.
We further look into the stability of the counterfactual methods in terms of sufficiency (Table P) and the explanation size (Table Q) across different Gnn architectures. The sufficiency results (Table P) show large variations produced by the same method on the same dataset due to the different architectures and message passing schemes. For instance, RCExplainer produces sufficiency of .10 and .93 on the AIDS dataset for GAT and GIN, respectively. In terms of explanation size(Table Q), RCExplainer is stable against different Gnn architectures. However, consistent with previous stability results, Cf is more unstable than RCExplainer and the worst stability is shown by Clear.
Stability to feature noise: Table R presents the impact of feature noise on counterfactual explanations in Mutag and Mutagenecity. In both datasets, the features are binary in nature. We sample of nodes uniformly at random and the flip a randomly sampled feature in this node. We then vary and measure how stability reacts to the injected noise. We observe that Cf and Clear are markedly more stable than RCExplainer in Mutag. This outcome is not surprising, considering that RCExplainer exclusively addresses topological perturbations, while both Cf and Clear accommodate perturbations encompassing both topology and features. In Mutaganecity, RCExplainer exhibits slightly higher stability than Cf.
@@
| Mutagenicity | 1.01 ±0.19 | 1.36 ±1.46 | 1.23 ±1.23 | 1.0 ±0.0 | 2.78 ±0.98 | 3.50 ±1.73 | 6.18 ±3.20 | 3.27 ±1.51 | OOM | OOM | OOM | OOM |
| Proteins | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | NA | NA | 3.40 ±1.48 | 1.92 ±0.71 | OOM | OOM | OOM | OOM |
| Mutag | 1.1 ±0.22 | 1.0 ±0.0 | NA | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 19.38 ±1.03 | 2.0 ±0.0 | 17.15 ±1.62 | 16.5 ±1.44 | NA | 19.11 ±3.28 |
| IMDB-B | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 8.57 ±4.99 | NA | 6.0 ±0.0 | 10.0 ±0.0 | 218.62 ±0.0 | 343.02 ±49.16 | 344.32 ±86.03 | 247.51 ±56.72 |
| AIDS | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 5.25 ±0.35 | 3.60 ±1.57 | 2.0 ±0.0 | 2.25 ±0.64 | 164.95 ±47.93 | 577.46 ±88.55 | 161.34 ±40.94 | 136.68 ±29.34 |
| ogbg-molhiv | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 10.45 ±4.43 | 7.58 ±4.09 | 21.92 ±11.60 | 5.46 ±2.17 | OOM | OOM | OOM | OOM |
C Necessity and Reproducibility for the Factual Explainers: Results for Sec. 4.3
@@
| 0 (no noise) | 0.4 ±0.06 | 1.01 ±0.19 | 1.0 ±0.0 | 0.50 ±0.05 | 2.78 ±0.98 | 1.0 ±0.0 |
| 10 | 0.43 ±0.04 | 1.01 ±0.19 | 0.96 ±0.04 | 0.49 ±0.04 | 2.11 ±1.06 | 0.92 ±0.06 |
| 20 | 0.47 ±0.06 | 1.0 ±0.0 | 0.95 ±0.04 | 0.53 ±0.06 | 1.73 ±0.94 | 0.86 ±0.08 |
| 30 | 0.50 ±0.04 | 1.0 ±0.0 | 0.94 ±0.04 | 0.61 ±0.04 | 1.68 ±0.91 | 0.86 ±0.09 |
| 40 | 0.52 ±0.05 | 1.0 ±0.0 | 0.93 ±0.06 | 0.54 ±0.05 | 1.47 ±0.73 | 0.87 ±0.09 |
| 50 | 0.49 ±0.05 | 1.0 ±0.0 | 0.93 ±0.06 | 0.67 ±0.01 | 1.54 ±1.06 | 0.86 ±0.08 |
We have defined Necessity (inferring on the residual graph obtained after removing the explanation) and (2) Reproducibility [26] (retraining the Gnn on a subset of residual graphs and then making predictions on the remaining unseen residual graphs) in Sec. 4.3. We present the results here. In both cases, we evaluate performance using sufficiency which in this context, measures the ratio of graphs for which the Gnn prediction on the residual graph is the same as in the original graph. After removing the explanation, we expect sufficiency to be low.
The results are presented in Figs. I and J. While sufficiency is indeed low in necessity (Fig. I), this is not the case in reproducibility (Fig. J). These findings suggest that while current factual explainers effectively explain the model, they do not provide a comprehensive explanation of the underlying data. The fact that the Gnn can regain prediction accuracy when retrained indicates the presence of other signals that the initial factual explanation failed to capture.
Counterfactual methods. Please note that we have omitted the results for counterfactual methods in Section 4.3. This is because the necessity of counterfactuals is defined as 1-sufficiency, making it a derived metric. To achieve reproducibility, it is necessary to retrain the Gnn; however, the label set for counterfactual graphs differs from the one used to train the original black-box Gnn. Consequently, comparing their performance becomes impractical.
D Visualization of Explanations
@@
In Figs. K and L, we engage in a visual analysis of the explanations provided by various Gnn explainers. The graphs presented in these figures represent mutagenic molecules sourced from the Mutag dataset. Several insights emerge from this analysis.
@@
Factual: The mutagenic attribute of the molecule in Fig. K stems from the presence of the NO group attached to the benzene ring[59, 14]. As a result, the optimal explanation entails pinpointing this specific benzene ring in conjunction with the NO group. Notably, we observe that while certain explainers identify fragments of this subgraph, with Cf achieving the highest overlap, many also highlight bonds originating from regions outside the authentic explanatory context. Adding to the intrigue, the explanation offered by RCExplainer stands out due to its compactness, resulting in commendable statistical performance. However, this succinct explanation lacks meaning in the eyes of a domain expert. Consequently, a pressing need arises for real-world datasets endowed with ground truth explanations, a resource that the current field unfortunately lacks.
@@
Counterfactuals: Fig. L illustrates two molecules, with Molecule 1 (top row) being identical to the one shown in Fig. K. The optimal explanation involves eliminating the NO component, a task accomplished solely by Cf in the case of Molecule 1. While the remaining explanation methods can indeed alter the Gnn prediction by implementing the changes described in Figure L, two critical insights emerge. First, statistically, RCExplainer is considered a better explanation than Cf since its size is compared to 3 of Cf. However, our interaction with multiple chemists clearly indicated their preference towards Cf since eliminates the entire NO group. Second, chemically infeasible explanations are common as evident from Clear for both molecules and Cf in molecule 2. Both fail to adhere to valency rules, a behavior also noted in Sec. 4.4.
E Additional Experiments on Sparsity Metric
@@
Counterfactual explainers: Sparsity is defined as the proportion of edges from the graph that is retained in the counter-factual [61]; a value close to is desired. In node classification, we compute this proportion for edges in the , i.e., the -hop neighbourhood of the target node . We present the results on sparsity metric on node and graph classification tasks in Table S and T, respectively. For node classification, we see CF-GnnExplainer continues to outperform (Table S). The results are consistent with our earlier results in Table 5. Similarly, Table T shows that RCExplainer continues to outperform in the case of graph classification (earlier results show a similar trend in Table 4).
@@
Factual explainers: For experiments with factual explainers, we report results of the necessity metric on varying degrees of sparsity (Recall Fig. I). Note that size acts as a proxy for sparsity in this case. This is because sparsity only involves the normalized size where it is normalized by the number of total edges. So sparsity can be obtained by normalizing all perturbation sizes in the plots to get the sparsity metric. For counterfactual explainers, we do not supply explanation size as a parameter. Hence, computing the sparsity of the predicted counterfactual becomes relevant.
@@
F TAGExplainer Variants
TAGExplainer has two stages. We define TAGExplainer (1) as when we apply only the first stage and get explanations, whereas TAGExplainer (2) applies both stages. Figure M compares performance of these two variants.
G Factual Explainers on OGBG-Molhiv
Figure O demonstrates a new dataset OGBG-Molhiv for three factual explainers. On this dataset, three factual methods are close to each other in terms of sufficiency performance.

### v3

@@ A.7 Feasibility
Factual explanations: The feasibility metric is commonly used in the context of counterfactual graph explainers because it measures how feasible it is to achieve a specific counterfactual outcome. In other words, it assesses the likelihood of a counterfactual scenario being realized given the constraints and assumptions of the underlying base model. On the other hand, factual explainers aim to explain why a model makes a certain prediction based on the actual input data. They do not involve any counterfactual scenarios, so the feasibility metric is not relevant in this context. Instead, factual explainers may use other metrics such as sufficiency and reproducibility to provide insights into how the model is making its predictions. Therefore, we have not used feasibility metrics for factual explanations.
B Sufficiency of Factual Explanations under Topological Noise
We check the sufficiency of the factual explanations under noise for four different datasets. Figure F demonstrates the results, including when there is no noise (i.e., when X = 0). We set the explanation size (i.e., the number of edges) to 10 units. We observe that, in most cases, increasing noise results in a decrease in the sufficiency metric for the Mutagenicity and AIDS datasets, which is expected. However, for the Proteins and IMDB-B datasets, even though there are still drops for some methods, others remain stable in sufficiency across different noise levels. This demonstrates that, despite the changes in explanations caused by noise, Gnn may still predict the same class under noisy conditions.
C Stability
In addition to stability against topological noise, different seeds, and different Gnn architecture, we also analyze stability against feature pertubation and stability against topological adversarial attack.
For feature perturbation, we first select the percentage of nodes to be perturbed (). Then, perturbation operation varies depending on the nature of node features (continuous or discrete). For the Proteins dataset (continuous features); for each feature , we compute its standard deviation . Then we sample a value uniformly at random from . The feature value is perturbed to . For other datasets (discrete features), for each selected node, we flip its feature to a randomly sampled feature.
For topology adversarial attack, we follow the flip edge method from Wan et al. (2021) with a query size of one and vary the number flip count across datasets.
C.1 Factual Explainers
Stability against feature perturbation:
Figure G illustrates the outcomes, demonstrating a continuation of the previously observed trends. Among these trends, we see that there is one clear winner for both datasets. However, PGExplainer performs better than most datasets in both datasets (with discrete and continuous features). On the other hand, the transductive method GNNExplainer performs very poorly in both datasets compared to other methods (i.e., inductive), which further provides evidence that transductive methods are poor in stability for factual explanations.
@@ C.1 Factual Explainers
Adversarial attack on topology: Figure H demonstrates the performance of four factual methods on these evasion attacks for four datasets. The behavior of the factual methods is similar to the topological noise attack explained in Section 4.2 and the feature perturbations results. When an adversarial attack is compared to random perturbations (Fig. 3), we observe higher deterioration in stability, which is expected since adversarial edge flip attack aims every possible edge in the graph rather than only considering nonexistent edges. Similar to feature perturbation, GNNExplainer (transductive) is affected more by the adversarial attack.
C.2 Counterfactual Explainers
@@
Figure I presents the average Jaccard similarity results, indicating the similarity between the counterfactual graph predicted as an explanation for the original graph and the noisy graphs at varying levels of perturbations. Additionally, Figure J demonstrates the performance of different explainers in terms of sufficiency and size as the degree of noise increases. This provides insights into how these explainers handle higher levels of noise.
RCExplainer outperforms other baselines by a significant margin in terms of size and sufficiency across datasets, as shown in Fig. J. However, the Jaccard similarity between RCExplainer and CF for counterfactual graphs is nearly identical, as shown in Fig. I. CF benefits from its transductive training on noisy graphs. CLEAR’s results are not shown for Proteins and Mutagenicity datasets due to scalability issues. In the case of IMDB-B dataset, CLEAR is highly unstable in predicting counterfactual graphs, indicated by a low Jaccard index (Fig. I). Additionally, CLEAR demonstrates high sufficiency but requires a large number of edits, indicating difficulty in finding minimal-edit counterfactuals (Fig. J).
Overall, RCExplainer seems to be the model of choice when topological noise is introduced, and it is significantly faster than Cf because it is inductive. Further, it is better than CLEAR as the latter does not scale for larger datasets and is inferior in terms of sufficiency and size as well.
Stability against explainer instances: Table M provides an overview of the stability exhibited among explainer instances trained using three distinct seeds. Notably, we observe a substantial Jaccard index, indicating favorable stability, in the case of RCExplainer and Cf explainers. Conversely, Clear fails to demonstrate comparable stability. These findings align with the outcomes derived from Table 5. Specifically, when RCExplainer and Cf are successful in identifying a counterfactual, the resultant counterfactual graphs are obtained through a small number of perturbations. Consequently, the counterfactual graphs exhibit similarities to the original graph, rendering them akin to one another. However, this trend does not hold for Clear, as it necessitates a significantly greater number of perturbations.
| RCExplainer | CF | CLEAR | |||||||
| Dataset / Seeds | |||||||||
| Mutagenicity | ± | ± | ± | ± | ± | ± | OOM | OOM | OOM |
| Proteins | ± | ± | ± | NA | NA | NA | OOM | OOM | OOM |
| Mutag | ± | ± | ± | ± | ± | ± | ± | ± | ± |
| IMDB-B | ± | ± | ± | ± | ± | ± | ± | ± | ± |
| AIDS | ± | ± | ± | NA | ± | NA | ± | ± | ± |
| ogbg-molhiv | ± | ± | ± | ± | ± | ± | OOM | OOM | OOM |
Similarly, as additional results, we also present the variations in terms of sufficiency (Table N) and the explanation size (Table O) produced by the methods using three different seeds. In terms of sufficiency, the methods show stability while varying the seeds. However, the results are drastically different for explanation size. Table O present the results. We observe that RCExplainer is consistently the most stable method while Cf is worse. The worst stability is shown by Clear and this observation is consistent with the previous results.
@@
Stability against Gnn architectures: Table P shows the stability of the explainers across different Gnn architectures. Similar to our factual setting (Table 8), we assess the stability by computing the Jaccard coefficient between the explained predictions of the indicated Gnn architecture and the default Gcn model. Unsurprisingly, the stability of the explainer highly depends on the dataset. RCExplainer is also the most stable among all the explainers, and the produced high values indicate that the method is agnostic towards the variations in different message aggregating schemes of the architectures.
We further look into the stability of the counterfactual methods in terms of sufficiency (Table Q) and the explanation size (Table R) across different Gnn architectures. The sufficiency results (Table Q) show large variations produced by the same method on the same dataset due to the different architectures and message passing schemes. For instance, RCExplainer produces sufficiency of .10 and .93 on the AIDS dataset for GAT and GIN, respectively. In terms of explanation size(Table R), RCExplainer is stable against different Gnn architectures. However, consistent with previous stability results, Cf is more unstable than RCExplainer and the worst stability is shown by Clear.
@@
| Mutagenicity | OOM | OOM | OOM | OOM | ||||||||
| Proteins | OOM | OOM | OOM | OOM | ||||||||
| Mutag | NA | NA | NA | |||||||||
| IMDB-B | NA | |||||||||||
| AIDS | NA | |||||||||||
Stability to feature noise: Table S presents the impact of feature noise on counterfactual explanations in Mutag and Mutagenicity. We observe that Cf and Clear are markedly more stable than RCExplainer in Mutag. This outcome is not surprising, considering that RCExplainer exclusively addresses topological perturbations, while both Cf and Clear accommodate perturbations encompassing both topology and features. In Mutaganecity, RCExplainer exhibits slightly higher stability than Cf.
@@
| 0 (no noise) | ||||||
| 10 | ||||||
| 20 | ||||||
| 30 | ||||||
| 40 | ||||||
| 50 | ||||||
D Necessity: Results for Sec. 4.3
For necessity, we remove explanations from graphs and measure the ratio of graphs for which the Gnn prediction on the residual graph is flipped. We expect that removing more explanations will lead to more necessity; however, we do not necessarily expect necessity to be higher because our factual explainers are not trained to make the residual graph counterfactual. Fig. K presents the necessity performance on six datasets for all factual methods with varying explanation sizes from 1 to 10. The trend aligns with our expectations, as the removal of more explanations increases the necessity. The value of necessity varies between datasets. Proteins and IMDB-B datasets have graphs with larger sizes (in terms of the number of edges), which aligns with the small necessity score. On the other hand, datasets with relatively smaller graphs have higher necessity scores. Note that our factual methods do not optimize residual graphs to be counterfactuals; this might be another reason for the low values.
E Reproducibility: Results for Sec. 4.3
Reproducibility can be measured two different ways. (1) Retraining using only explanation graphs called Reproducibility, (2) retraining using only residual graphs called Reproducibility. Both metrics is a ratio of an Gnn accuracy compared to the original Gnn accuracy. We provide the math definitions in Table 3. In our figures, we separated SubgraphX to an independent table, because we could only obtain explanations of test graphs for SubgraphX (refer to our disclaimer in Sec. A.4)
| Size (#edges in the explanation) | ||||||||||
| Datasets | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 |
| Mutagenicity | ||||||||||
| Mutag | ||||||||||
| IMDB-B | ||||||||||
| AIDS | ||||||||||
| NCI1 | ||||||||||
Figure L and Table T illustrate the Reproducibility performance of seven factual methods against the size of explanations for six datasets. Reproducibility increases with more edges in the explanations for the most cases, as expected. However, reaching a score of 1.0 is challenging even when selecting the most crucial edges. This suggests that explanations do not capture the full picture for Gnn predictions.
Figure M and Table U illustrate the Reproducibility performance of seven factual methods against the size of explanations for six datasets. Reproducibility remains high even when the most crucial edges from the graphs are removed. This demonstrates that the explainers hardly capture the real cause of the Gnn predictions.
| Size (#edges in the explanation) | ||||||||||
| Datasets | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 |
| Mutagenicity | ||||||||||
| Mutag | ||||||||||
| IMDB-B | ||||||||||
| AIDS | ||||||||||
| NCI1 | ||||||||||
F Visualization of Explanations
@@
In Figs. N and O, we engage in a visual analysis of the explanations provided by various Gnn explainers. The graphs presented in these figures represent mutagenic molecules sourced from the Mutag dataset. Several insights emerge from this analysis.
@@
Factual: The mutagenic attribute of the molecule in Fig. N stems from the presence of the NO group attached to the benzene ringYing et al. (2019b); Debnath et al. (1991). As a result, the optimal explanation entails pinpointing this specific benzene ring in conjunction with the NO group. Notably, we observe that while certain explainers identify fragments of this subgraph, with Cf achieving the highest overlap, many also highlight bonds originating from regions outside the authentic explanatory context. Adding to the intrigue, the explanation offered by RCExplainer stands out due to its compactness, resulting in commendable statistical performance. However, this succinct explanation lacks meaning in the eyes of a domain expert. Consequently, a pressing need arises for real-world datasets endowed with ground truth explanations, a resource that the current field unfortunately lacks.
@@
Counterfactuals: Fig. O illustrates two molecules, with Molecule 1 (top row) being identical to the one shown in Fig. N. The optimal explanation involves eliminating the NO component, a task accomplished solely by Cf in the case of Molecule 1. While the remaining explanation methods can indeed alter the Gnn prediction by implementing the changes described in Figure O, two critical insights emerge. First, statistically, RCExplainer is considered a better explanation than Cf since its size is compared to 3 of Cf. However, our interaction with multiple chemists clearly indicated their preference towards Cf since eliminates the entire NO group. Second, chemically infeasible explanations are common as evident from Clear for both molecules and Cf in molecule 2. Both fail to adhere to valency rules, a behavior also noted in Sec. 4.4.
G Additional Experiments on Sparsity Metric
@@
Counterfactual explainers: Sparsity is defined as the proportion of edges from the graph that is retained in the counter-factual Yuan et al. (2022); a value close to is desired. In node classification, we compute this proportion for edges in the , i.e., the -hop neighbourhood of the target node . We present the results on sparsity metric on node and graph classification tasks in Table V and W, respectively. For node classification, we see CF-GnnExplainer continues to outperform (Table V). The results are consistent with our earlier results in Table 6. Similarly, Table W shows that RCExplainer continues to outperform in the case of graph classification (earlier results show a similar trend in Table 5).
@@
Factual explainers: For experiments with factual explainers, we report results of the necessity metric on varying degrees of sparsity (Recall Fig. K). Note that size acts as a proxy for sparsity in this case. This is because sparsity only involves the normalized size where it is normalized by the number of total edges. So sparsity can be obtained by normalizing all perturbation sizes in the plots to get the sparsity metric. For counterfactual explainers, we do not supply explanation size as a parameter. Hence, computing the sparsity of the predicted counterfactual becomes relevant.
@@
H TAGExplainer Variants
TAGExplainer has two stages. We define TAGExplainer (1) as when we apply only the first stage and get explanations, whereas TAGExplainer (2) applies both stages. Figure P compares performance of these two variants.
I Factual Explainers on OGBG-Molhiv
Figure R demonstrates a new dataset OGBG-Molhiv for three factual explainers. On this dataset, three factual methods are close to each other in terms of sufficiency performance.
J Existing Benchmarking Studies on Gnn Explainability
GraphFrameX Amara et al. (2022) and GraphXAI Agarwal et al. (2023) represent two notable benchmarking studies. While both investigations have contributed valuable insights into Gnn explainers, certain unresolved investigative aspects persist.
Inclusion of counterfactual explainability: GraphFrameX and GraphXAI have focused on factual explainers for GNNs. Prado-Romero et al. (2023) has discussed methods and challenges, but benchmarking on counterfactual explainers remains underexplored.
K Recommendations in practice
The choice of the explainer depends on various factors and we make the following recommendations.
For counter-factual reasoning, we recommend RCExplainer for graph classification and CF-GNNExplainer for node classification.
For factual reasoning, if the goal is to do node or graph classification, we need to first decide if we need an inductive reasoner or transductive. While an inductive reasoner is more suitable we want to generalize to large volumes of unseen graphs, transductive is suitable for explaining single instances. In case of inductive, we recommend RCExplainer, while for transductive GNNExplainer stands out as the method of choice. These two algorithms had the highest sufficiency on average. In addition, RCExplainer displayed stability in the face of noise injection.
For different task generalization beyond node and graph classification, one can consider using TAGExplainer.
In high-stakes applications, we recommend RCExplainer due to consistent results across different runs and robustness to noise.
For both types of explainers, the transductive methods are slow. So, if the dataset is large, it is always better to use an inductive explainer over transductive ones.
While the above is a guideline, we emphasize that there is no one-size-fits-all benchmark for selecting the ideal explainer. The choice depends on the characteristics of the application in hand and/or the dataset. The above flowchart takes these factors into account to streamline the decision process. We have now added a flowchart to help the user in selecting the most appropriate. This recommendation has now also been added as a flowchart (Fig. S).